OpenStudy (anonymous):

find the area of the triangle whose sides have the given lengths. a=1, b=2, c=2

OpenStudy (amistre64):

use herons formula, or find an angle with the law of cosine

OpenStudy (anonymous):

@ Chiomatn use the method i gave you it is valid for all

OpenStudy (anonymous):

is s=2.5?

OpenStudy (anonymous):

@Chiomatn93 , I just told you the answer few minutes back. The heron's formula

OpenStudy (amistre64):

 c^2 = a^2 + b^2 -2ab cos(C) c^2-a^2-b^2 ------------ = cos(C) -2ab c^2-a^2-b^2 cos^-1 (------------) = C -2ab 2^2 sinC -------- = area 2

OpenStudy (anonymous):

that wasnt the answer... u just gave me the formula

OpenStudy (anonymous):

@Chiomatn93 , I am not here to give answers :(

OpenStudy (anonymous):

If you want some help , I can help you

OpenStudy (anonymous):

all i want is the steps and answer.. i already know the formula.. is s=2.5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@ chiomatn93 be confident in what you are you would make mistakes I also did just keep on trying

OpenStudy (anonymous):

@Chiomatn93, you need a calculator, not us :P

OpenStudy (anonymous):

Let \[s =\frac {a+b+c}{2}=\frac 5 2\\ Area= \sqrt{ s(s-a)(s-b)(s-c)}=\frac {\sqrt{15}}{4} \]

OpenStudy (amistre64):

if never been fond of the heron formula :)

OpenStudy (anonymous):

im getting decimals

OpenStudy (anonymous):

\[ \frac{\sqrt{15}}{4}\approx 0.968246 \]